OpenStudy (bohotness):

If the radius of the circle above is 5 cm, what is the circumference of the circle in terms of ? A. 10 cm B. 20 cm C. 5 cm D. 25 cm

OpenStudy (bohotness):

@Data_LG2

OpenStudy (girlover):

2piR=C

OpenStudy (anonymous):

Remember pi = 3.14

OpenStudy (texaschic101):

C = 2 * pi * r C = 2 * 3.14 * 5 C = wait...in terms of what ?

OpenStudy (texaschic101):

this problem does not make sense....is the problem written correctly ? It would end up being : 10 * pi...or 10 * 3.14

OpenStudy (bohotness):

so 10 times 3.13?

OpenStudy (bohotness):

i mean 3.14

OpenStudy (bohotness):

no If the radius of the circle above is 5 cm, what is the circumference of the circle in terms of ? A. 10 cm B. 20 cm C. 5 cm D. 25 cm